# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ V-A3 Attribute Inference Attacks
Attribute inference attacks aim to deduce the presence of specific attributes or characteristics of data samples within the training data of LM agents. For instance, such attacks can be exploited to infer the proportion of images with a specific artist style in the training data of a text-to-image agent, potentially leading to privacy breaches for providers of these training images. Pan et al. [192] systematically investigate the privacy risks associated with attribute inference attacks in LLMs. Through four diverse case studies, they validate the existence of sensitive attribute inference (e.g., identity, genome, healthcare, and location) within the training data of general-purpose LLMs. Wang et al. [193] explore the property existence inference attack against generative models, aiming to determine whether any samples with a target property are contained in the training data. Their study shows that most generative models are susceptible to property existence inference attacks, and they validate this vulnerability in stable diffusion models.
@@ V-A4 Countermeasures to LM Memorization Risks in LM Agent Networks
Existing countermeasures to mitigate memorization risks of LM agents primarily focus on data pre-processing during pre-training and fine-tuning phases. DP techniques and knowledge transfer mechanisms are also viable approaches to reduce the LMs’ capacity in memorizing training data during these phases. Additionally, it is a common strategy to detect and validate privacy disclosure risks of LM agents before deployment.
  • Data sanitization can effectively mitigate memorization risks by identifying and excluding sensitive information from training data. By replacing sensitive information with meaningless symbols or synthetic data, and removing duplicated sequences, it is possible to defend against privacy threats that exploit the memorization characteristics of LM agents. Kandpal et al. [194] demonstrate that the rate at which memorized training sequences are regenerated is superlinearly related to the frequency of those sequences in the training data. Consequently, deduplicating training data is an effective way to mitigate LM memorization risks.
  • Adding DP noises to training data and model gradients during pre-training and fine-tuning phases can effectively mitigate privacy leakages due to LM memorization. For instance, Hoory et al. [195] propose a novel differentially private word-piece algorithm, which achieves a trade-off between model performance and privacy preservation capability.
  • Knowledge distillation [113] has been widely adopted as an intuitive technique to preserve privacy. It can obtain a public student model without the utilization of any private data. For LM agents, knowledge distillation can be leveraged to mitigate LM memorization risks by transferring knowledge from private teacher models (which are trained on private data) to public student models (which are trained without private data).
  • Privacy leakage detection and assessment can effectively mitigate LM memorization risks before deploying an LM agent for practical services, thereby enabling service providers to modify the model based on assessment results. For instance, Kim et al. [196] propose ProPILE, an innovative probing tool to evaluate privacy intrusions in LLMs. ProPILE can be employed by service providers to evaluate the levels of PII leakage of their LLM agents.
@@ V-B LM Stealing & Prompt Stealing Attacks to LM Agent Networks
In LM agent networks, LM stealing risks (including LM’s parameters, hyperparameters, and specific training processes) and prompt stealing risks (prompts are considered as commodities to generate outputs) are two types of intellectual property (IP)-related privacy threats to LM agents, as illustrated in Fig. 31. LM-related information may inherently contain private information, and smart attackers can further infer private data from the extracted information through carefully crafted privacy threats. Besides, prompts typically contain user inputs that not only indicate user intent, and requirements, but may also involve confidential business logic.
@@ V-B1 Model Stealing Attacks
In model stealing attacks, adversaries aim to extract model information, such as model parameters or hyperparameters, by querying models and observing the corresponding responses, subsequently stealing target models without access the original data [197]. Recently, Krishna et al. [198] demonstrate that language models (e.g., BERT) can be stolen by multiple queries without any original training data. Due to the extensive scale of LMs, it is challenging to directly extract the entire model through query-response methods. Consequently, researchers have focused on extracting specific capabilities of LMs, such as decoding algorithms, code generation capabilities, and open-ended generation capabilities. Naseh et al. [199] demonstrate that adversaries can steal the type and hyperparameters of an LM’s decoding algorithms at a low cost through query APIs. Li et al. [200] investigate the feasibility and effectiveness of model stealing attacks on LLMs to extract the specialized code abilities. Jiang et al. [201] propose a novel model stealing attack, which leverages the adversarial distillation to extract knowledge of ChatGPT to a student model through a mere 70k training data, and the student model can achieve comparable open-ended generation capabilities to ChatGPT.
@@ V-B2 Prompt Stealing Attacks
With the advancement of LM agent services, high-quality prompts designed to generate expected content have acquired substantial commercial value. These prompts can be traded on various prompt marketplaces, such as PromptSea [^12] https://www.promptsea.io/ and PromptBase [^13] https://promptbase.com/. Consequently, a new privacy threat called prompt stealing attack has emerged, where an adversary aims to infer the original prompt from the generated content. This attack is analogous to the model inversion attack in traditional AI, which involves reconstructing the input based on the output of an AI model [202]. Shen et al.[203] conduct the first study on prompt stealing attack in text-to-image generation models, and propose an effective prompt stealing attack method named PromptStealer. PromptStealer utilizes a subject generator to infer the subject and a modifier detector to identify the modifiers within the generated image. Sha et al.[204] extend the prompt stealing attack to LLMs, using a parameter extractor to determine the properties of original prompts and a prompt reconstructor to generate reversed prompts. Hui et al.[205] propose an innovative closed-box prompt extraction framework, called PLEAK, which employs incremental search techniques and post-processing methods to optimize adversarial queries for the purpose of extracting system prompts from LLM applications. The effectiveness of PLEAK is demonstrated through its evaluation on real-world applications hosted on the Poe platform  [^14] https://poe.com.
TABLE VIII: Summary of Key Literature on Privacy Threat Countermeasures to LM Agent Networks
@@ Methods
$\star$ Reduce the privacy leakage risk due to memorization of training data in language models
@@
$\dagger$ AUROC(Area under ROC curve), TPR
@@ V-B3 Countermeasures to Model & Prompt Stealing Attacks in LM Agent Networks
Existing countermeasures to model and prompt stealing attacks involve both IP verification (e.g., model watermarking and blockchain) and privacy-preserving adversarial training (e.g., adversarial perturbations), as detailed below.
  • Model watermarking is an innovative technique to protect IP rights and ensure accountability for LM agents. By embedding watermarks to target LMs, the ownership of LMs can be authenticated by verifying the watermarks, thereby preventing unauthorized use or infringement. For instance, Kirchenbauer et al. [206] propose a watermarking algorithm utilizing a randomized set of “green” tokens during the text generation process, where the model watermark is verified by a statistical test with interpretable p-values.
  • Blockchain can be employed as a transparent platform to verify IP rights due to its inherent immutability and traceability [207]. The owner of LMs can record the develop logs, version information, and hash values of LM parameters on blockchain, ensuring the authenticity and completeness of the recorded information. Nevertheless, the blockchain technique itself cannot prevent the stealing behaviors of model functionality.
  • Adversarial examples is an effective method to prevent prompt-stealing attacks while maintaining the quality of the generated content by adding optimized perturbations to generated content. Shen et al. [203] propose an intuitive defense mechanism named PromptShield, which employs the adversarial example technique to add a negligible perturbation on generated images, thereby defending against the proposed prompt stealing attack PromptStealer. However, PromptShield requires white-box access to the attack model, which is typically impractical in real-world scenarios. Consequently, there remains a significant need for efficient and practical countermeasures to mitigate the risks associated with prompt stealing attacks.
@@ V-C Other Privacy Threats to LM Agent Networks
  • Sensitive query attacks: In LM agent services, the LM may memorize sensitive personal or organizational information in user queries, resulting in potential privacy leakages. For instance, Samsung employees leveraged ChatGPT for code auditing without processing the confidential information in Apr. 2023, inadvertently exposing the company’s commercial secrets including source code of the new program.
  • Privacy leakage in multi-agent interactions: LM agent services typically necessitate seamless collaboration among multiple LM agents to address complex user queries, where each agent is tasked with solving particular sub-problems of the queries. Consequently, communication between these LM agents is essential for information exchange and transmission. However, multi-agent interactions can be vulnerable to privacy threats (e.g., eavesdropping, compromised agent, and man-in-the-middle attacks), leading to potential user privacy breaches. Since interactions in LM agent services typically occur through natural language and semantic meanings, traditional methods such as homomorphic encryption and secure multi-party computation struggle to effectively safeguard the privacy of these semantically rich interactions. It remains a challenge to design new strategies tailored to these specific vulnerabilities to preserve privacy during multi-agent interactions.
@@ V-D Summary and Lessons Learned
There are primarily two types of privacy threats to LM agents: LM memorization threats, and IP-related privacy threats. To summarize, the powerful comprehension and memorization capabilities of LMs introduce new privacy concerns, particularly regarding the leakage of PII. Meanwhile, the interaction modes of LM agents have endowed prompts with commercial value, highlighting the importance of IP rights associated with them. Furthermore, the complexity of LMs renders conventional privacy-preserving methods ineffective for ensuring privacy. Therefore, to comprehensively safeguard privacy within LM agent systems, researchers should develop effective and innovative privacy protection techniques tailed for connected LM agents. Additionally, it is imperative for governments and authoritative organizations to advance legislation process related to privacy breaches and IP of LM agent services.
VI Future Research Directions

### v3

@@ V-A3 Attribute Inference Attacks
Attribute inference attacks aim to infer the presence of specific attributes or characteristics of data samples within the training data of LM agents. For instance, such attacks can be exploited to infer the proportion of images with a specific artist style in the training data of a text-to-image agent, potentially leading to privacy breaches for the owners of these images. Pan et al. [173] conduct a comprehensive examination of privacy vulnerabilities posed by attribute inference attacks in LLMs. Through four diverse case studies, they show the potential for inferring sensitive attributes (e.g., identity, genome, healthcare, and location information) in the training data of general-purpose LLMs. Besides, Wang et al. [174] study the property existence inference attack against generative models, aiming to detect the presence of specific data samples associated with a target property in the training data. Their study reveals that most generative models, including stable diffusion models, are susceptible to property existence inference attacks.
@@ V-A4 Countermeasures to LM Memorization Risks in LM Agent Networks
Existing countermeasures to mitigate memorization risks of LM agents primarily focus on data pre-processing during pre-training and fine-tuning phases. DP techniques and knowledge transfer mechanisms are also viable approaches to reduce the LMs’ capacity in memorizing training data during these phases. Additionally, detecting and assessing privacy risks before deploying LM agents is a common approach.
  • Data sanitization: Data sanitization can effectively mitigate memorization risks by identifying and excluding sensitive information from training data. This can be achieved by replacing sensitive information with meaningless symbols or synthetic data, and removing duplicated sequences. Kandpal et al. [175] demonstrate the superlinear correlation between the regeneration frequency of training sequences and their original occurrence rates in the original training dataset. Therefore, deduplication of training data is an effective way to mitigate LM memorization risks.
  • Adding DP noises to training data and model gradients: Applying DP noises to the training data and model gradients during pre-training and fine-tuning phases can mitigate privacy leakage due to LM memorization. For instance, Hoory et al. [176] propose a differentially private word-piece algorithm that establishes a trade-off between model performance and privacy preservation capability.
  • Knowledge distillation: Knowledge distillation [54] is a widely adopted technique to preserve privacy. It involves transferring knowledge from private teacher models (trained on sensitive data) to public student models (trained without access to private data). This approach allows the student model to retain valuable knowledge without memorizing specific details from the private training data, thereby mitigating memorization risks.
  • Privacy leakage detection and assessment: Detecting and assessing privacy leakage before deploying an LM agent for practical services can help mitigate memorization risks, allowing LM agents to make necessary adjustments to the model based on the assessment results. For instance, Kim et al. [177] propose ProPILE, a probing tool to evaluate privacy intrusions in LLMs. ProPILE enables service providers to evaluate the levels of PII leakage in their LLM agents.
@@ V-B LM Stealing & Prompt Stealing Attacks to LM Agent Networks
In LM agent networks, LM stealing risks (including the theft of LM parameters, hyperparameters, and specific training processes) and prompt stealing risks (prompts are considered as commodities to generate high-quality outputs) are two types of intellectual property (IP)-related privacy threats to LM agents, as illustrated in Fig. 34. LM-related information may inherently contain private information, and smart attackers can infer private data from the extracted information through carefully crafted privacy threats. Besides, prompts typically contain user inputs that not only indicate user intent and requirements, but may also encapsulate confidential business logic, representing another potential vulnerability.
@@ V-B1 Model Stealing Attacks
In model stealing attacks, adversaries aim to extract model information, such as LM parameters and hyperparameters, by querying models and observing their responses, subsequently stealing target models without access the original data [178]. Krishna et al. [179] show that language models (e.g., BERT) can be stolen via multiple queries without any original training data. Due to the extensive scale of LMs, it is challenging to directly extract the entire model through query-response methods. Thus, researchers focus on stealing specific capabilities of LMs, such as decoding, code generation, and open-ended generation abilities. Naseh et al. [180] demonstrate that adversaries can steal the type and hyperparameters of an LM’s decoding algorithms through query APIs at a low cost. Li et al. [181] explore the feasibility of stealing LLMs’ specialized code abilities. Additionally, Jiang et al. [182] present a new model stealing attack, which leverages adversarial distillation to extract knowledge from ChatGPT into a student model through only 70k training samples, enabling the student model to achieve comparable open-ended generation capabilities to ChatGPT.
@@ V-B2 Prompt Stealing Attacks
High-quality prompts are essential for LM agent applications, such as task orchestration prompts for complex tasks. These prompts, often proprietary, hold commercial value and are traded on platforms such as PromptSea [^4] https://www.promptsea.io/ and PromptBase [^5] https://promptbase.com/. Consequently, prompt stealing attacks have emerged, where adversaries aim to infer the original prompt from the generated content. Analogous to model inversion attacks, these attacks aim to reconstruct the original prompt from generated content [183]. Shen et al.[184] pioneer PromptStealer, the first attack on text-to-image models, capable of extracting subject and modifier information from generated images. Sha et al.[185] extend prompt stealing attacks to LLM agents, using a parameter extractor to infer original prompt attributes and a prompt reconstructor to recreate them. Hui et al.[186] propose PLEAK, a closed-box prompt extraction framework that employs incremental search and adversarial query optimization to extract prompts from LM-based applications, demonstrating its effectiveness on real-world services hosted on the Poe platform [^6] https://poe.com.
TABLE IX: Summary of Key Literature on Privacy Threats & Countermeasures to LM Agent Networks
@@ Methods
$\star$ Reduce the privacy leakage risk due to language model memorization of training data
@@
$\dagger$ Area under ROC curve (AUROC), TPR
@@ V-B3 Countermeasures to Model & Prompt Stealing Attacks in LM Agent Networks
Existing countermeasures to model and prompt stealing attacks involve both IP verification (e.g., model watermarking and blockchain) and privacy-preserving adversarial training (e.g., adversarial perturbations), as outlined below.
  • Model watermarking techniques embed watermarks into LMs to protect IP rights, verify ownership, and ensure accountability for LM agents. Specifically, the ownership of LMs can be authenticated by verifying the embeded watermarks, thereby preventing unauthorized use or infringement. For instance, Kirchenbauer et al. [187] propose a watermarking algorithm utilizing a randomized set of “green” tokens during text generation, where the model watermark is verified by a statistical test with interpretable p-values.
  • Blockchain technology can enforce transparency, immutability, and traceability for verifying IP rights [188]. The owner of LMs can record develop logs, version information, and hash values of LM parameters on the blockchain, ensuring authenticity. Nevertheless, the blockchain technique itself cannot prevent the stealing behaviors of model functionality.
  • Adversarial examples can defend against prompt stealing attacks while preserving output quality, by adding optimized perturbations to the generated content. For instance, Shen et al. [184] propose an intuitive defense mechanism named PromptShield, which employs the adversarial example technique to add negligible perturbations on generated images, thereby defending against their proposed prompt stealing attack PromptStealer. However, PromptShield requires white-box access to the attack model, which may be impractical in real-world scenarios. Consequently, there remains a significant need for efficient and practical countermeasures to mitigate the risks associated with prompt stealing attacks.
@@ V-C Other Privacy Threats to LM Agent Networks
  • Sensitive query attacks: In LM agent services, LM agents may inadvertently memorize sensitive personal or organizational information from user queries, leading to potential privacy leaks. For instance, Samsung employees used ChatGPT for code auditing without adequately sanitizing confidential information in Apr. 2023, inadvertently exposing the company’s commercial secrets including source code of the new program [189].
  • Privacy leakage in multi-agent interactions: LM agent networks rely on seamless collaboration, necessitating frequent communication for information exchange in addressing complex tasks. However, multi-agent interactions can be vulnerable to privacy threats [190], including eavesdropping, compromised agents, and man-in-the-middle attacks, causing potential privacy breaches. Since interactions between LM agents typically occur through natural language and semantic meanings, traditional security mechanisms such as secure multi-party computation and homomorphic encryption struggle to effectively safeguard the privacy of these semantically rich interactions.
@@ V-D Summary and Lessons Learned
There are primarily two categories of privacy threats to LM agents: LM memorization threats and IP-related privacy threats. To summarize, the powerful comprehension and memorization abilities of LMs raise new privacy concerns, particularly regarding the leakage of PII. Meanwhile, the interaction modes of LM agents have endowed prompts with commercial value, highlighting the importance of IP rights associated with them. Furthermore, the complexity of LMs renders conventional privacy-preserving methods less effective in ensuring comprehensive privacy. Therefore, to comprehensively protect privacy within LM agent systems, researchers should develop effective and innovative privacy protection techniques tailed for connected LM agents. Additionally, it is imperative for governments and regulatory bodies to advance legislation addressing privacy breaches and IP protection in LM agent services. Table IX summarizes existing typical research on privacy threats and countermeasures to LM agent networks.
Figure 35: The taxonomy of reliability issues in LM agent networks.
VI Reliability Issues & Countermeasures to Large Model Agent Networks
Figure 36: Category of hallucinations in the context of LM agents.
Decision-making reliability is crucial for practically deploying LM agent services. Particularly, the hallucination phenomenon in LM agents refers to the generation of erroneous or illogical outputs that deviate from user inputs, generated context, or real-world conditions, posing a significant risk to service reliability. In multi-agent scenarios, hallucinations can be exacerbated through agent interactions, amplifying and propagating errors across the system. Fig. 36 illustrates the taxonomy of reliability issues to LM agent networks. According to [28], we categorize the hallucination within the context of LM agents from the following four perspectives, as illustrated in Fig. 36.
VI-A Causes of Hallucinations of LM Agent Networks
TABLE X: Summary of Key Literature on Reliability Issues & Countermeasures to LM Agent Networks
Ref.
Reliability
Issues
$\star$ Purpose
$\bullet$ Advantages
$\circ$ Limitations
$\dagger$ Evaluation Metrics
Defense
Methods
[191]
Hallucinations in
table-to-text tasks
$\circ$ Annotation complexity, and lack of scalability
$\dagger$ BLEU and PARENT scores
Data
sanitization
[192]
Hallucinations of
generated tasks
$\dagger$ AUC-PR, Pearson correlation coefficient, and Spearman’s rank correlation coefficient
Hallucination
detection
[193]
Hallucinations of
QA datasets
$\dagger$ Accuracy, latency, and cost
RAG
[194]
Hallucinations of
realistic abstractive
summarization tasks
$\star$ Reduce hallucinations by assessing in synthesis tasks
$\circ$ Synthetic task design dependency, and model dependency
Instruction
tuning
VI-A1 Hallucination from Data and Knowledge
The primary cause of hallucinations in LM agents stems from the biased or imperfect nature of the training data and knowledge used for content generation, as below.
  • Annotation irrelevance: Irrelevant annotations in the collected dataset can lead to hallucinations in LM agents. For LLMs, source-reference divergence occurs due to heuristic data collection, implying that the selected sources may lack sufficient or accurate information for content generation [195]. Similarly, in LVMs, a substantial portion of instructions is synthesized using LLMs, which may not accurately correspond to the content depicted in the images, as generative models can be unreliable [196].
  • Distribution imbalance: The presence of distribution imbalance in the training data can exacerbate hallucinations. For instance, an overrepresentation of positive samples compared to negative ones in binary discrimination tasks can lead to models responding in a biased way, such as frequently answering “yes” when faced with a decision [196, 197].
  • Data duplication: If duplicates are not properly filtered during pre-processing, LMs may generate responses that closely resemble the patterns in those duplicates [198].
  • Hallucinatory knowledge: The training data and knowledge used in LM agents can also be inherently hallucinatory due to bias, outdated content, or fabrication [199]. LMs agents trained on such data are susceptible to replicate or even amplify hallucinations. Anomalous retrieval from the knowledge base further introduces biases into the dependent knowledge, thereby exacerbating the emergence of hallucinations in the generated content.
VI-A2 Hallucination from Training and Inference
Hallucinations in LM agents can occur not only due to biases in training data but also as a result of imperfections in the training and inference processes. Parikh et al. [191] demonstrate that hallucinations can arise even when the training dataset is nearly unbiased, with several factors contributing to the phenomenon.
  • Defective encoder and decoder components: Hallucinations can be caused by flaws in the encoder and decoder components of LM agents. The flawed representation learning in LM training exacerbates the unpredictability of the generation process, increasing the likelihood of hallucinated content [191]. Moreover, the sampling-based top-p decoding strategy, which is inherently random, has been shown to correlate with an increase in hallucinated content [200].
  • Parametric knowledge: LMs store knowledge from training data in their model parameters, which can improve performance on downstream tasks [123]. However, recent studies suggest that LMs tend to prioritize parametric knowledge over user input during content generation [201], implying that biased parametric knowledge can result in a large amount of hallucinated information.
  • Alignment process: Hallucinations can also result from a flawed alignment process. If the necessary prior knowledge is not adequately acquired during LM pre-training phase, the agent may produce hallucinated responses. Additionally, sycophancy, where the agent generates answers that align more with the user’s viewpoint than with accurate and credible information, can contribute to hallucinations [202].
  • Token generation and hallucination snowballing: The token generation process itself can contribute to hallucinations through a phenomenon known as hallucination snowballing. This occurs when the LM agent persists in early errors for the sake of self-consistency, rather than correcting them, which can compound the hallucination in later generated content [203].
VI-A3 Hallucination from Multi-agent Interactions
Multi-agent communication introduces new hallucination risks, exacerbated by conflicting agent outputs and misinformation propagation.
  • Conflicting agents. Due to the diverse objectives, strategies, and knowledge bases among LM agents, conflicting viewpoints or misinformation between multiple agents can inadvertently lead to contradictory responses and influence the final output. Adversarial manipulation may intensify these inconsistencies, leading to more severe hallucinations.
  • Cascading hallucinations. In multi-agent systems, a single agent’s hallucination can trigger a cascading effect, where misinformation from one agent is accepted and propagated by others across the network [37, 64], particularly in vertical collaboration paradigms. Addressing hallucinations in LM agent networks necessitates both individual agent-level correction mechanisms and robust inter-agent validation mechanisms to prevent hallucination propagation.
VI-B Countermeasures to Reliability Issues in LM Agent Networks
Existing countermeasures to mitigate hallucinations in LM agents focus on the following aspects.
  • Data sanitization: Filtering unreliable data from training datasets is a fundamental approach to reduce hallucinations. For instance, the manually-revised ToTTo dataset is validated to improve factuality in table-to-text generation tasks [191], while RefinedWeb, an automatically filtered and deduplicated web dataset, can mitigate hallucinations and enhance LLM reliability [199].
  • Reinforcement learning: By integrating external feedback mechanisms and reward-based constraints, reinforcement learning discourages LM agents from generating hallucinatory information. This approach enables agents to recognize their limitations and avoid answering beyond their knowledge scope instead of fabricating untruthful responses [204, 79]. For instance, GPT-4 collects user-annotated unfactual data and generates synthetic closed-domain hallucinated synthetic data to train a reward model, thereby reducing hallucination tendencies.
  • Hallucination detection: Identifying hallucinations post-generation allows for sample regeneration to eliminate inaccuracies. Techniques such as SelfcheckGPT [192] and INSIDE [205] assess consistency across multiple generated responses to detect hallucinated content.
  • Truthful external knowledge & RAG: Truthful RAG techniques can help anchor LM-generated content in factual information [53, 79]. By incorporating verified external knowledge sources and established contextual knowledge, LM agents can reduce hallucinations in generated responses [193].
  • Instruction tuning: Optimizing instruction sets or fine-tuning LMs on robust instruction datasets reduces hallucinations. For instance, SYNTRA [194] evaluates hallucinations on synthesis tasks and refines instructions to reduce hallucinations on downstream tasks. Besides, Liu et al. [196] mitigate hallucinations in LVMs by constructing a robust instruction dataset named LRV-Instuction and fine-tuning LVMs on it.
  • Post-processing: Refining LM-generated content through post-processing can correct hallucinations in LM responses [206, 207]. For instance, RARR [206] utilizes search engines to gather relevant external knowledge and refine responses, while LURE [207] reduces hallucinations by correcting object-level inaccuracies during post-processing.
  • Model architecture optimization: Structural modifications to LMs can reduce hallucinations. Multi-branch decoder [208] and uncertainty-aware decoder [209] are two examples of modified decoder structures to mitigate hallucinations.
  • Meta programming: Meta-programming frameworks can guide and promote collaboration among LM agents, thereby reducing hallucinations in complex tasks [64].
VI-C Summary and Lessons Learned
The reliability of LM agents is primarily compromised by hallucinations, which arise from biased training data, imperfect learning and inference processes, and multi-agent interactions. While countermeasures such as data sanitization, hallucination detection, and RAG can mitigate these issues, the probabilistic nature of generative models makes hallucinations inherently unavoidable. Future research should continuously focus on hallucination mitigation methods, including multi-agent consensus mechanisms and dynamic credibility audit frameworks, to enhance the reliability of both generated content and decision-making processes in LM agent networks. Table X summarizes existing typical researches on reliability issues and countermeasures to LM agent networks.
VII Future Research Directions